Solving 2 equations: x-3y=4
(Trying to find x & y) 2x+y=6

Answer by scott8148(6628)

many methods for solution

adding 3y to first equation ___ x = 3y + 4

substituting into second equation ___ 2(3y + 4) + y = 6 ___ 7y + 8 = 6 ___ y = -2/7

substituting ___ x = 3(-2/7) + 4 = 22/7




multiplying first equation by 2 ___ 2x - 6y = 8

subtracting second equation ___ -7y = 2 ___ y = -2/7

substituting ___ x - 3(-2/7) = 4 ___ x = 22/7